A spinner is divided into quarters: one blue, one red, one yellow, and one green. a 6 sided die is also rolled. What is the likelihood that a person spins a blue and rolls a even number.
Answer #1

Ans:

Probability of landing on blue color=1/4

Probability of landing on even number=2/6=1/3

(i.e. getting a 2,4 or 6)

Probability that a person spins a blue and rolls a even number=(1/4)*(1/3)=1/12=0.0833